THE ν_{2} BANDS OF BrNO_{2} (NITRYL BROMIDE) AROUND 787 CM^{−1}

F. K. TCHANA, \underline{J. ORPHAL}, I. KLEINER, CNRS, Laboratoire de Photophysique Moléculaire, 91405 Orsay Cedex, France; B. REDLICH, Institut für Anorganische Chemie, Universität Münster, 48149 Münster, Germany; D. SCHEFFLER, H. WILLNER, Institut für Anorganische Chemie, Universität Duisburg, 47048 Duisburg, Germany; R. MBIAKE, Centre de Physique Atomique Moléculaire et Optique Quantique, Université de Douala, Cameroun; O. BOUBA, Université de Ngaoundéré, Cameroun.

Nitryl bromide, BrNO_{2}, is produced in the marine troposphere by heterogeneous reactions between nitrogen oxides and bromine containing aerosols.^{a} Its peak concentrations reach values of several 10^{7} cm^{−3}.^{b} In polluted coastal areas, therefore, BrNO_{2} is an important species in the release of bromine from the ocean into the atmosphere.^{c} The ν_{2} fundamental bands of ^{79}BrNO_{2} and ^{81}BrNO_{2}, located around 787 cm^{−1} (12.7 μm), were recorded using a highresolution Fourier-transform infrared spectrometer. A total of nearly 5000 transitions with J ≤ 80 and K_{a} ≤ 30 were reproduced using a Watson-type A-reduced Hamiltonian with a root-mean-square deviation of better than 5 × 10^{−4}cm^{−1}. Rotational and centrifugal distortion constants for the v_{2} states have been determined, as well as an improved set of ground state constants for both isotopomers. Due to their sharp Q branches falling into an atmospheric window, the ν_{2} bands might be useful for future attempts to detect atmospheric BrNO_{2}.

aW. Behnke, V. Sheer, and C. Zetzsch, J. Aerosol. Sci. 25, S277-S278, 1994.

bJ. Moldanova and E. Ljungström, J. Geophys. Res. D106, 1271-1296, 2001.

cR. Vogt, P. J. Crutzen, and R. Sander, Nature 383, 327-330, 1996.